• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:49:49 on July30,2003

    Entry number 108204

    keyword=day shift summary

    Shift Summary:

    Shift worker: R. Roche, R. Subedi, J. P. Chen

    8:00 continue pressure curve for 3He at E=2.134 GeV, P=2.125 (elastic)
    run 3042, 7:53-8:04, He3 pressure=90 psig, PS1=4, PS2=2, 1.94M events
    run3043, 8:07-8:17, He3 pressure=140 psig, PS1=4, PS2=2, 2.37M
    8:30 Pump out He3, Pressure curve for N2
    run 3044, 8:31-8:41, N2 pressure 140 psig, PS1=35, PS2=20, 2M
    run 3045, 8:42-8:52, N2 pressure 90 psig, PS1=35, PS2=20, 1.47 M
    run 3046 short, stop to change prescalers
    run 3047, 8:54-~9:00, N2 pressure=35 psig, PS1=12, PS2=5, DAQ problem, no endrun
    9:00 right DAQ problem: TS0 not responding, endrun failed, reset TS0, kcoda
    all is fine.
    run 3048, 9:06-9:17 , N2 pressure=35 psig, PS1=12, PS2=5, 2.33M

    9:20 change to kin 5.10, P=2.005 GeV (forgot to change Septum current, left at
    196.73A, should be 185.62A)!
    NMR=41.13%, EPR=40.13%
    9:45 all set, asking for beam back.
    9:50 kin 5.10, 0 degree, BHWP out
    run 3049, 9:52-10:10, 3M
    run 3050, 10:12-10:20, 1.5M
    rotating target to 180, target half wave plate in.
    kin 5.10, BHWP out, target 180 degree.
    run 3051, 10:38-10:53, 3M
    run 3052, 10:53-11:07, 3M

    11:07 changing to kin 5.11, p=1.871 GeV, Septum=173.22A,
    target 0 degree, BHWP out.
    Bleedthrough measurement=2%
    11:35 Magnets are set
    run 3053, short, PS too large.
    run 3054, 11:36-11:50, 3M
    run 3055, 11:50-12:04, 3M
    run 3056, 12:05-12:20, 3M
    12:20 asked to have BHWP in.
    run 3057, 12:23-12:36, 3M
    run 3058, 12:36-12:51, 3M
    run 3059, 12:51-13:05, 3M
    13:06 rotated to 270 degree.
    Beam off due to RF problem.
    14:00 Beam study to 16:00, then beam energy and pass change.
    Hall A in controlled access.